OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = 2/3x − 10 y = 2/3x + 10 y = −3/2x − 10 y = −3/2x + 10

OpenStudy (anonymous):

First, put the equation in the problem into the form y = mx + b to get the slope (m). A perpendicular line will have slope of -1/m. The "new slope value" Now, use equation\[y - y _{1} = (new slope value)(x - x _{1})\] and then rearrange into the same form of y = mx + b.

OpenStudy (anonymous):

In the last equation in my post above, the x1 and y1 are specific coordinates that correspond to the point "through which the line passes" which is given to you in the problem which is (-8, 2). "x" and "y" without the subscript are variables for the equation you are determining.